# Pistis-RAG: A Scalable Cascading Framework Towards Trustworthy Retrieval-Augmented Generation

### v1

@@ Abstract.
In Greek mythology, Pistis symbolized good faith, trust, and reliability, echoing the core principles of RAG in LLM systems. Pistis-RAG, a scalable multi-stage framework, effectively addresses the challenges of large-scale retrieval-augmented generation (RAG). Each stage plays a distinct role: matching refines the search space, pre-ranking prioritizes semantically relevant documents, and ranking aligns with the large language model’s (LLM) preferences. The reasoning and aggregating stage supports the implementation of complex chain-of-thought (CoT) methods within this cascading structure.
We argue that the lack of strong alignment between LLMs and the external knowledge ranking methods used in RAG tasks is relevant to the reliance on the model-centric paradigm in RAG frameworks. A content-centric approach would prioritize seamless integration between the LLMs and external information sources, optimizing the content transformation process for each specific task.
Critically, our ranking stage deviates from traditional RAG approaches by recognizing that semantic relevance alone may not directly translate to improved generation. This is due to the sensitivity of the few-shot prompt order, as highlighted in prior work (lu2021fantastically, ). Current RAG frameworks fail to account for this crucial factor.
We introduce a novel ranking stage specifically designed for RAG systems. It adheres to information retrieval principles while considering the unique business scenario captured by LLM preferences and user feedback. Our approach integrates in-context learning (ICL) methods and reasoning steps to incorporate user feedback, ensuring efficient alignment.
Experiments on the MMLU benchmark demonstrate a 9.3% performance improvement. The model and code will be open-sourced on GitHub. Experiments on real-world, large-scale data validate our framework’s scalability.
@@ 1. Introduction
Figure 1. Brief of RAG Framework.
@@
In recent years, LLMs have demonstrated remarkable capabilities in natural language understanding and generation tasks. These models, trained on vast datasets representing a wide array of human knowledge, excel at generating coherent and contextually relevant text across various domains  (achiam2023gpt, ). Despite their advancements, LLMs are inherently limited by the scope and quality of their training data. They cannot often incorporate real-time or domain-specific knowledge not present in their training corpus, leading to challenges in accuracy and relevance, particularly in dynamic environments.
To mitigate these limitations, RAG has emerged as a promising paradigm. RAG enhances LLMs by augmenting their input with additional information retrieved from external sources at inference time. This approach not only enriches the context of LLM-generated responses but also enables them to leverage up-to-date and domain-specific knowledge without the need for extensive retraining [cite needed]. By integrating ICL from retrieved knowledge, LLMs can adapt and perform better on tasks requiring specialized knowledge or current events.
Current RAG frameworks face significant challenges that hinder their widespread adoption. These challenges primarily lie in two areas: accurately retrieving relevant knowledge and striking a balance between generation quality and computational efficiency. Additionally, integrating retrieved information with LLM-generated outputs remains a non-trivial task, requiring careful consideration of factors like semantic alignment, prompt engineering, and user preferences (as discussed in  (yu2023augmentation, )). Traditional RAG approaches may also overlook the nuances of how LLMs process and prioritize information, leading to suboptimal performance in real-world applications.
@@ 3.2.6. Stage 6: Aggregating (Consistency Checking)
Figure 5. The Multi-Stage Cascading Retrieval and Generation (Pistis-RAG) Framework is designed to optimize information retrieval and response generation for LLMs. It comprises five stages: Matching, Ranking, Reranking, Reasoning, and Aggregating. In the Matching stage, retrieval algorithms select pertinent documents from repositories. Semantic analysis in the Ranking stage refines document scores based on query relevance. Reranking aligns document ranking with LLM preferences for coherent responses. Reasoning enhances response diversity, exploring multiple sequences. The Aggregation stage synthesizes responses for coherence. Pistis-RAG employs advanced techniques such as Cross-Encoder methods and Concurrent Chain-of-Thought generation to enhance performance and response quality. It is a comprehensive solution for complex queries and contextually relevant responses.
@@ 3.3. Architecture
The Reasoning Service Enhanced Reasoning with Parallel Inference and Expert Routing. The Reasoning Service leverages an advanced Chain of Thought to execute parallel inference across LLMs, generating diverse results for subsequent aggregation. This service takes prompts and retrieved information as input, producing a series of reasoned outputs. By harnessing LLMs, it extracts insights, draws conclusions, and answers inquiries based on the provided data. The outcome is a comprehensive set of reasoned responses that comprehensively address the user’s intent.
@@
Merging Concurrent Inference Results: This service seamlessly combines outcomes from concurrent inference processes, guaranteeing coherence and consistency. Self-consistency checks and similar techniques are employed to validate and harmonize the aggregated results.
@@
  • Citation and Transparency: Credibility is bolstered by incorporating citations to trustworthy sources within industry contexts. This may involve referencing established sources and providing transparency regarding data origin. Additionally, the reasoning process, such as the Chain of Thought and decision-making steps, can be showcased to provide deeper insights.
@@ 4. Ranking Details
In this section, we outline the ranking process, differentiating it from the commonly discussed concept of re-rankers in RAG discourse. The ranking stage in information retrieval systems plays a crucial role in aligning content delivery with business objectives and user preferences. The ranking stage transcends this, encompassing a broader range of tasks essential for optimizing the retrieval and presentation of information.
@@ 4.2. Ranking for LLM Prompt Order Sensitivity
Existing rule-based studies, such as those focusing on cue word placement, suggest that placing examples at the beginning or end may be more effective. However, we assume that this preference cannot be determined by rules alone; rather, it requires a statistic learning model to capture this preference distribution accurately.
The RAG system addresses prompt order sensitivity by employing a ranking mechanism. This mechanism ensures that the most relevant and informative prompt examples are presented to the LLM first. This alignment between the ranking and the ideal order of prompts leads to the following:
@@
4.3. Ranking Stage Problem Definition
The ranking problem in online systems involves improving the order of items presented to users based on various factors beyond just relevance. This can be achieved by leveraging user feedback to understand their preferences for the content.
User feedback can be obtained through actions like copying the content (indicating strong preference), regenerating (suggesting mild preference for an alternative), or disliking the content (clear disapproval). By analyzing these feedback labels, we aim to learn a ranking model that optimizes the order of items within a list to match user preferences better.
Formally, let’s denote our dataset as $\mathcal{D}$, containing a set of ordered few-shot examples $\{(p_{i},x_{i},y_{i})\}$ and corresponding user feedback vectors $L$. Here, $p_{i}$ represents the user intent, $x_{i}$ represents the input example, $y_{i}$ represents the desired output (potentially the originally ranked list), and $L_{i}$ is a vector containing the specific feedback labels (copying, regenerating, disliking) associated with the $i$-th pair.
The RAG ranking stages’ objective is to develop a ranking model that utilizes these feedback signals to optimize the ranking sequences produced by the large language model, ultimately enhancing the user experience and satisfaction with the system.
@@ 4.4. Listwide Labels
This part delves into the complexities of learning from implicit user feedback in Listwide Learning to Rank (LTR) methods. Specifically, we focus on scenarios where explicit relevancy labels for all items in a list are absent.
Our primary focus is on leveraging indirect signals derived from user feedback, which reflect how the ranking of content impacts the results produced by LLMs. These indirect signals are termed Listwide Labels.
@@ 4.5. Ranking Model Alignment
Additionally, we integrate techniques for managing user feedback dynamics and evolving user preferences over time, ensuring our ranking model remains adaptive and responsive to changes in user behavior.
@@ 5. EXPERIMENTS ON PUBLIC DATASETS
Our experimental analysis delves into prominent public datasets MMLU(hendrycks2020measuring, ), which comprises a substantial volume of data, with human annotators assigning explicit answers to each question in the selection.
However, our primary focus lies in learning from feedback labels, which are more abundant and reflective of real-world scenarios. Hence, we simulate feedback labels to mirror realistic user behavior observed in various industries. This simulation approach is detailed in the subsequent sections.
@@ 5.1.1. Collecting Few-shot Examples
Within our system architecture, we seamlessly integrate highly rated user feedback into our contextual learning example database for RAG utilization. Specifically, we incorporate precise question-answer pairs sourced from public datasets as supplementary ICL examples $D$. To ensure impartial evaluation, we filter out answers identical to questions during retrieval (though in the live online system, such cases are directly cached for future retrieval). This meticulous approach establishes a simulated example of learning from an external knowledge base, enriching the system’s performance through augmented retrieval data.
@@ 5.1.2. Simulating User Behavior
We refine user behavior data from open-source datasets to simulate user feedback through the following steps:
@@
  • No Answer ($L(y)=Negative$): Represents outputs lacking an answer and belonging to the set of non-responses $Y_{\text{no\_answer}}$.
This formalization illustrates a continuous learning system wherein feedback from real user interactions refines and enhances the model, aligning its outputs more closely with user expectations and real-world applications. Specifically, correct answers ($Y_{\text{correct}}$) resemble text copying, incorrect answers ($Y_{\text{incorrect}}$) resemble regeneration, and no answers ($Y_{\text{no\_answer}}$) correspond to negative user feedback.
@@ 5.2. Datasets
In assessing the effectiveness of the Pistis-RAG framework, it is essential to leverage diverse datasets that offer a comprehensive evaluation of its capabilities. Two prominent datasets, MMLU and NetEval, serve as invaluable resources in this regard, providing insights into the framework’s adaptability and performance across various domains.
The MMLU dataset is a valuable resource for researchers who are developing speech recognition and natural language processing systems. The dataset has been used to train various speech recognition systems, and it has been shown to improve the performance of these systems. The dataset has also been used to develop natural language processing systems that can understand and generate spoken language.
@@ 5.3. Experimental Setup
Retrieval: We used BEG-M3 to retrieve the top 10 candidate few-shots based on Milvus. BEG-M3 is a dense vector similarity search engine that can efficiently retrieve similar vectors from a large collection of vectors. Milvus is a vector database that stores and manages vectors.
@@ 6. Conclusions
We revisited the Retrieval-Augmented Generation problem through a content-centric lens, uncovering potential shortcomings in large-scale deployments. Firstly, a mismatch exists between the intended and actual function of the "re-ranker" component, suggesting it acts more like a pre-ranker. We addressed this by proposing distinct ranking and re-ranking stages within RAG, along with a listwide ranking approach. Furthermore, we introduced reasoning and aggregation stages specific to AIGC services, paving the way for a more comprehensive and high-performing system.

### v2

# Pistis-RAG: A Scalable Cascading Framework Towards Content-Centric Retrieval-Augmented Generation

@@ Abstract.
In Greek mythology, Pistis symbolized good faith, trust, and reliability. Drawing inspiration from these principles, Pistis-RAG is a scalable multi-stage framework designed to address the challenges of large-scale retrieval-augmented generation (RAG) systems. This framework consists of distinct stages: matching, pre-ranking, ranking, reasoning, and aggregating. Each stage contributes to narrowing the search space, prioritizing semantically relevant documents, aligning with the large language model’s (LLM) preferences, supporting complex chain-of-thought (CoT) methods, and combining information from multiple sources.
Our ranking stage introduces a significant innovation by recognizing that semantic relevance alone may not lead to improved generation quality, due to the sensitivity of the few-shot prompt order, as noted in previous research (lu2021fantastically, ). This critical aspect is often overlooked in current RAG frameworks.
We argue that the alignment issue between LLMs and external knowledge ranking methods is tied to the model-centric paradigm dominant in RAG systems. We propose a content-centric approach, emphasizing seamless integration between LLMs and external information sources to optimize content transformation for specific tasks.
Our novel ranking stage is designed specifically for RAG systems, incorporating principles of information retrieval while considering the unique business scenarios reflected in LLM preferences and user feedback. We simulated feedback signals on the MMLU benchmark, resulting in a 9.3% performance improvement. Our model and code will be open-sourced on GitHub. Additionally, experiments on real-world, large-scale data validate the scalability of our framework.
@@ 1. Introduction
Figure 1. Brief of RAG Framework: from the initial bi-encoder encoding to mixed retrieval, and cross-encoder re-ranking.
[…]
Diagram showing the RAG framework stages: initial bi-encoder encoding, hybrid retrieval, and re-ranking.
@@
[…]
Diagram comparing traditional Information Retrieval (IR) systems and the RAG system. The IR systems include ranking and re-ranking processes, while the RAG system does not involve these steps, highlighting a mismatch in re-rankers.
In recent years, LLMs have demonstrated remarkable capabilities in natural language understanding and generation tasks. These models, trained in vast datasets that represent a wide range of human knowledge, excel in generating coherent and contextually relevant text in various domains  (achiam2023gpt, ). Despite their advancements, LLMs are inherently limited by the scope and quality of their training data. They cannot often incorporate real-time or domain-specific knowledge not present in their training corpus, leading to challenges in accuracy and relevance, particularly in dynamic environments.
To mitigate these limitations, RAG  (lewis2020retrieval, ) has emerged as a promising paradigm. RAG enhances LLMs by augmenting their input with additional information retrieved from external sources at inference time. This approach not only enriches the context of LLM-generated responses but also enables them to leverage up-to-date and domain-specific knowledge without the need for extensive retraining. By integrating ICL from the retrieved knowledge, LLMs can adapt and perform better on tasks that require specialized knowledge or current events.
Current RAG frameworks face significant challenges that hinder their widespread adoption. These challenges primarily lie in two areas: accurately retrieving relevant knowledge and striking a balance between generation quality and computational efficiency. Additionally, integrating retrieved information with LLM-generated outputs remains a nontrivial task, requiring careful consideration of factors like semantic alignment, prompt engineering, and user preferences (as discussed in  (yu2023augmentation, )). Traditional RAG approaches may also overlook the nuances of how LLMs process and prioritize information, leading to suboptimal performance in real-world applications.
@@ 3. The Pistis-RAG Framework
[…]
Diagram of a Content-Centric Cascading System (Pistis-RAG) illustrating the process of content retrieval and generation from a content-centric perspective, incorporating external data sources.
@@
[…]
Diagram showing a model-centric perspective from a LangChain-like framework, contrasting with a content-centric cascading system. Emphasis is on the model’s role in processing tasks, focusing on the model’s capabilities rather than just content flow and transformation.
@@ 3.2.6. Stage 6: Aggregating (Consistency Checking)
Figure 5. The Multi-Stage Cascading Retrieval and Pistis-RAG Framework is designed to optimize information retrieval and response generation for LLMs. It comprises five stages: Matching, Pre-Ranking, Ranking, Reasoning, and Aggregating. In the Matching stage, retrieval algorithms select pertinent documents from repositories. Semantic analysis in the Pre-Ranking stage refines document scores based on query relevance. Ranking aligns document ranking with LLM preferences for coherent responses. Reasoning enhances response diversity, exploring multiple sequences. The Aggregation stage synthesizes responses for coherence. Pistis-RAG employs advanced techniques such as Cross-Encoder methods and Concurrent Chain-of-Thought generation to enhance performance and response quality. It is a comprehensive solution for complex queries and contextually relevant responses.
[…]
Diagram of the Pistis-RAG Framework, detailing its five stages: Matching, Ranking, Reranking, Reasoning, and Aggregating. The framework optimizes information retrieval and response generation for LLMs using advanced techniques like Cross-Encoder methods and Concurrent Chain-of-Thought generation.
@@ 3.3. Architecture
The Reasoning Service Enhanced Reasoning with Parallel Inference and Expert Routing. The Reasoning Service leverages an advanced Chain of Though  (wei2022chain, ) to execute parallel inference across LLMs, generating diverse results for subsequent aggregation. This service takes prompts and retrieved information as input, producing a series of reasoned outputs. By harnessing LLMs, it extracts insights, draws conclusions, and answers inquiries based on the provided data. The outcome is a comprehensive set of reasoned responses that comprehensively address the user’s intent.
@@
Merging Concurrent Inference Results: This service seamlessly combines outcomes from concurrent inference processes, guaranteeing coherence and consistency. Self-consistency  (wang2022self, ) checks and similar techniques are employed to validate and harmonize the aggregated results.
@@
  • Citation and Transparency: Credibility is bolstered by incorporating citations  (li2024citation, ) to trustworthy sources within industry contexts. This may involve referencing established sources and providing transparency regarding data origin. Additionally, the reasoning process, such as the Chain of Thought and decision-making steps, can be showcased to provide deeper insights.
@@ 4. Ranking Details
In this section, we outline the ranking process, differentiating it from the commonly discussed concept of rerankers in RAG discourse. The ranking stage in information retrieval systems plays a crucial role in aligning content delivery with business objectives and user preferences. The ranking stage transcends this, encompassing a broader range of tasks essential for optimizing the retrieval and presentation of information.
@@ 4.2. Ranking for LLM Prompt Order Sensitivity
Existing rule-based studies, such as those that focus on cue word placement, suggest that placing examples at the beginning or end may be more effective. However, we assume that this preference cannot be determined by rules alone; rather, it requires a statistic learning model to capture this preference distribution accurately.
The RAG system addresses prompt order sensitivity by employing a ranking mechanism. This mechanism ensures that the most relevant and informative prompt examples are presented to the LLM first. This alignment between the ranking and the ideal order of prompts leads to the following.
@@
4.3. Ranking Problem Definition
The ranking problem in online systems involves improving the order of items presented to users based on various factors beyond just relevance. This can be achieved by using user feedback to understand their preferences for the content.
User feedback can be obtained through actions such as copying the content (indicating a strong preference), regenerating (suggesting a mild preference for an alternative), or disliking the content (clear disapproval). By analyzing these feedback labels, we aim to learn a ranking model that optimizes the order of items within a list to better match user preferences.
Formally, let us denote our dataset as $\mathcal{D}$, containing a set of examples $\{(p_{i},x_{i},y_{i},f_{i})\}$, where:
  • $p_{i}$ represents the user’s intent.
  • $x_{i}$ represents the ordered few-shot example list.
  • $y_{i}$ represents the end-to-end generated output.
  • $f_{i}$ represents the specific feedback label (copying, regenerating, disliking) associated with the $i$-th pair.
The objective of the RAG ranking stages is to develop a ranking model that utilizes these feedback signals to optimize the ranking sequences produced by the large language model, ultimately enhancing the user experience and satisfaction with the system.
@@ 4.4. Listwide Labels
This part delves into the complexities of learning from implicit user feedback in Listwide Learning to Rank (LTR) methods. Specifically, we focus on scenarios where explicit relevancy labels are absent for all items in a list.
Our primary focus is on using indirect signals derived from user feedback, which reflect how the ranking of content impacts the results produced by LLMs. These indirect signals are termed Listwide Labels.
@@ 4.5. Ranking Model Alignment
In addition, we integrate techniques for managing user feedback dynamics and evolving user preferences over time, ensuring that our ranking model remains adaptive and responsive to changes in user behavior.
@@ 5. Experiments on Public Datasets
Our experimental analysis delves into the prominent public datasets MMLU(hendrycks2020measuring, ), which comprise a substantial volume of data, and human annotators assign explicit answers to each question in the selection.
However, our primary focus lies in learning from feedback labels, which are more abundant and reflective of real-world scenarios. Hence, we simulate feedback labels to mirror the realistic user behavior observed in various industries. This simulation approach is detailed in the subsequent sections.
@@ 5.1.1. Collecting Few-shot Examples
Within our system architecture, we seamlessly integrate highly rated user feedback into our contextual learning example database for the utilization of RAG. Specifically, we incorporate precise question-answer pairs sourced from public datasets as supplementary ICL examples $D$. To ensure impartial evaluation, we filter out answers identical to questions during retrieval (though in the live online system, such cases are directly cached for future retrieval). This meticulous approach establishes a simulated example of learning from an external knowledge base, enriching the system’s performance through augmented retrieval data.
@@ 5.1.2. Simulating User Behavior
We refine user behavior data from open source datasets to simulate user feedback through the following steps:
@@
  • No Answer ($L(y)=Negative$): Represents outputs lacking an answer and belonging to the set of nonresponses $Y_{\text{no\_answer}}$.
This formalization illustrates a continuous learning system in which feedback from real user interactions refines and enhances the model, aligning its output more closely with user expectations and real-world applications. Specifically, the correct answers ($Y_{\text{correct}}$) resemble text copying, the incorrect answers ($Y_{\text{incorrect}}$) resemble regeneration, and the lack of answers ($Y_{\text{no\_answer}}$) correspond to negative user feedback.
@@ 5.2. Datasets
In assessing the effectiveness of the Pistis-RAG framework, it is essential to leverage diverse datasets that offer a comprehensive evaluation of its capabilities. MMLU is an invaluable resource, providing insights into the framework’s adaptability and performance across various domains.
The MMLU dataset is a valuable resource for researchers developing speech recognition and natural language processing systems. The data set has been used to train various speech recognition systems and has been shown to improve the performance of these systems. The dataset has also been used to develop natural language processing systems that can understand and generate spoken language.
@@ 5.3. Experimental Setup
Matching: We used BEG-M3 to retrieve the top 10 candidate few-shots based on Milvus. BEG-M3 is a dense vector similarity search engine that can efficiently retrieve similar vectors from a large collection of vectors. Milvus is a vector database that stores and manages vectors.
@@ 6. Conclusions
We revisited the Retrieval-Augmented Generation problem through a content-centric lens, uncovering potential shortcomings in large-scale deployments. Firstly, a mismatch exists between the intended and actual function of the "re-ranker" component, suggesting it acts more like a pre-ranker. We addressed this by proposing distinct ranking and re-ranking stages within RAG and a listwide ranking approach. Furthermore, we introduced reasoning and aggregation stages specific to AIGC services, paving the way for a more comprehensive and high-performing system.